One positive number is one-fifth of another number. The difference between the two numbers is 188, find the numbers.
Answer by edjones(8007)

let b=biggest s=smallest
5s=b
b-s=188
replace b with 5s in the 2nd equation: 5s-s=188
4s=188
divide 4 into both sides: s=188/4
s=47
b=5*47=235
Check:
235-47=188
Right on!
Ed